consider an economy with only two industries , called Holi Apple company and Spencer juice company. Holi Apple company owns and operate apple groves . It sells some of its apples directly to the public. It sells the rest of its apples to Spencer juice company , which produces and sells apple juice. The following tables shows the transactions of each business during a year


TRANSACTIONS OF HOLI APPLE COMPANY

wages paid to workers $15,000

taxes paid to government $5,000

revenue received from sales of apples $35,000

apples sold to the public $10,000

apples sold to Spencer juice company $25,000

TRANSACTIONS OF SPENCER JUICE COMPANY

wages paid to workers of Spencer juice company $10,000

taxes paid to the government $2,000

apples bought from Holi apple company $25000

revenue received from sale of apple juice $40,000


using the data above , show that the product approach ,income approach and expenditure approach of national income accounting yield that same value of total economic activity(GDP)




1
Expert's answer
2021-06-22T15:32:52-0400

HOLI APPLE COMPANY

income approach

GDP = Total National Income + Sales Taxes + Depreciation + Net Foreign Factor Income

= 10000+25000

= $35000

expenditure approach

GDP = C + I + G +NX

= 35000+15000+5000

= $55000

SPENCER JUICE COMPANY

income approach

GDP = Total National Income + Sales Taxes + Depreciation + Net Foreign Factor Income

= 25000+40000

= $65000

expenditure approach

GDP = C + I + G +NX

= 10000+2000

= $12000